OpenStudy (anonymous):

x + y = 3 and y = x + 3

sam (.sam.):

?

sam (.sam.):

find x and y?

sam (.sam.):

x+y=3 x+(x+3)=3 2x+3=3 2x=0 x=0 y=3

OpenStudy (anonymous):

Use either elimination or substitution to find the solutions (if any)

sam (.sam.):

I use substitution

sam (.sam.):

x=0 y=3